I need help with my son's homework please. The problem is:
Consider the system {y=a+bx
y=16-7x} (Brackets are around both equations)
If this system has infinitely many solutions, then what must be the value of a?

Answer by rothauserc(4718)

Set the two equations = to each other, then
a + bx = 16 - 7x
a = 16 - bx - 7x
a = 16 - (b + 7)x